OpenStudy (senia01):

Transform this equation to slope-intercept form: 6x + 4y = 8. A. y equals begin fraction three over two end fraction times x plus four B. y equals begin fraction three over four end fraction times x plus two C. y equals negative begin fraction three over two end fraction times x plus two D. y equals negative begin fraction three over two end fraction times x plus four

OpenStudy (anonymous):

You just want to solve for y 6x+4y = 8; subtract 6x from each side 6x-6x+4y = 8-6x; simplify 4y = 8-6x; divide everything by 4 4y/4 = 8/4 -6x/4; simplify y = 2-6x/4; reduce and rearrange y = -3x/2+2 <-------Answer

OpenStudy (anonymous):

Correct Answer Is, Y=-x/2+2